A flat plate having dimensions 2m ×2m slides down an inclined plane at an angle of 1 rad to the horizontal at a speed of 6m/sec. The inclined plane is lubricated by a thin film of oil having a viscosity of 30 m pa.s. The plate has a uniform thickness of 20mm and a density of 40000km/m3. Determine the thickness of the lubricating oil film.

1
Expert's answer
2022-05-25T10:27:15-0400

"d=\frac{\eta\sin\alpha}{\rho lS}=1.2~mm."